# A Comunication Framework for Compositional Generation

### v1

@@ Abstract
Compositionality and compositional generalization—the ability to understand novel combinations of known concepts—are central characteristics of human language and are hypothesized to be essential for human cognition. In machine learning, the emergence of this property has been studied in a communication game setting, where independent agents (a sender and a receiver) converge to a shared encoding policy from a set of states to a space of discrete messages, where the receiver can correctly reconstruct the states observed by the sender using only the sender’s messages. The use of communication games in generation tasks is still largely unexplored, with recent methods for compositional generation focusing mainly on the use of supervised guidance (either through class labels or text). In this work, we take the first steps to fill this gap, and we present a self-supervised generative communication game-based framework for creating compositional encodings in learned representations from pre-trained encoder-decoder models. In an Iterated Learning (IL) protocol involving a sender and a receiver, we apply alternating pressures for compression and diversity of encoded discrete messages, so that the protocol converges to an efficient but unambiguous encoding. Approximate message entropy regularization is used to favor compositional encodings. Our framework is based on rigorous justifications and proofs of defining and balancing the concepts of Eficiency, Unambiguity and Non-Holisticity in encoding. We test our method on the compositional image dataset Shapes3D, demonstrating robust performance in both reconstruction and compositionality metrics, surpassing other tested discrete message frameworks.
Figure 1: General concept of the proposed framework. In a communication game a sender $S$ and receiver $R$ jointly aim to encode the observed latent states as symbolic messages. Useful length $\ell$ gives the minimum length of an initial segment of a $S$-encoded message, after which its $R$-reconstruction no longer improves. Above we show an example of images generated from increasing initial segments of a message, and in this case $\ell(x)=6$. Below we illustrate a typical alternation between interaction and imitation phases in IL: arrow tips indicate latent space $R$-encodings of increasing initial submessages of lenghts $1,\dots,\ell(x)$ of $S$-encoded messages.
Machine Learning, ICML
@@ 1 Introduction
Humans have the innate ability to segment and reimagine cognitive and perceptual representations, by modifying their components and characteristics, much like how we compose words into complex sentences to describe and understand the world around us (Arun, 2022).
One of the main challenges in modern Machine Learning is to faithfully capture this hability, also known as compositional generalization. In particular, the ability to imagine new combinations of known components translates directly into Compositionality in Generative models.
Compositional generation is directly relevant to the prominent challenges of learning long-tailed data distributions, in which models tend to mode collapse due to underrepresented classes (Saxena & Cao, 2021; Qin et al., 2023). It is often the case for long-tail distributions that several variable features of the data can be compositionally modified, producing a large combinatorial ensemble of different modes (Wiedemer et al., 2023), and thus compositional capabilities of a learning model will directly transfer to strong advances in Out Of Distribution (OOD) generalization (Liu et al., 2021; Abbasi et al., 2024).
The quest for efficient compositional priors for machine learning models has long been one of the main challenges of artificial intelligence. We should distinguish between several possible meanings for compositional generation:
  1. 1.
    Efficient generative transduction includes classical text-to-image generation tasks in which compositional dependencies are required for accurate OOD generation (Liu et al., 2022b; Leivada et al., 2023).
  2. 2.
    Controllable image generation requires to use, or detect and use, disentangled compositional features which will be controlled for correct generation (Nie et al., 2021; Casanova et al., 2023; Okawa et al., 2024)
  3. 3.
    Faithful generation of multimodal statistical distributions with latent compositional structure assumes that underlying multimodality of black-box target distributions, there is a compositional structure of the data. Here, no attempt at direct capture (or disentanglement) of this structure is made, and the focus is on imposing architecture priors to enhance the faithfulness of representation (Chowdhury et al., 2021; Latheef et al., 2024; Denamganai & Walker, 2020; Dessì et al., 2021).
We note that while the above three directions all require tackling the challenge of compositionality in different forms, cases 1 and 2 face the further difficulty of complex choice of metrics for testing their success. Here, we focus on direction 3 as it addresses a more fundamental question, in which generation accuracy objectively quantifies the success of capturing the full compositional complexity if we use a compositionally generated dataset. This relatively simpler setting allows us to focus precisely on the challenges of effective learning of compositional priors, leaving the extensions to more complex goals for future works.
Recent works have found a connection between simpler representations and compositionality (Elmoznino et al., 2024; Ren et al., 2023; Ren & Sutherland, 2024). Specifically, (Ren & Sutherland, 2024) found that compositional mappings are the simplest (smallest upper bound on Kolmogorov complexity) bijections between the underlying generating factors of a signal and the representation space.
Furthermore, as an independent benefit, from foundational works such as (Rissanen, 1978), penalizing the Kolmogorov Complexity (KC) of descriptions has been viewed as a computational version of Occam’s razor for learning models, inducing a kind of regularization leading to efficiency improvement for general learning frameworks (Hochreiter & Schmidhuber, 1997a; Hinton & Van Camp, 1993). In practice, indirect methods have to be adopted, because KC is impractical to compute, and furthermore in our case the coding underlying the compositional structure of our data is not assumed to be accessible. A promising approach has been the IL framework (Kirby, 2001), which implements active agents to learn a mapping from a meaning space into a signal space, modelled e.g. as a Definite Clause Grammar.
The IL mechanism can be modeled as a communication channel bottleneck (Havrylov & Titov, 2017; Kharitonov et al., 2020; Lerch & Sims, 2019; Resnick et al., 2019), with its success in KC regularization being justified via rate distortion theory (Tishby et al., 2000), but the main motivation for its validity comes from the analogy to the theory of how human and general languages are formed. Here the basic framework stems from sociolinguistic theory (Kirby & Hurford, 2002; Kirby et al., 2015; Blume et al., 1993; Bhui et al., 2021; Norris & Kalm, 2021; Rita et al., 2022; Lai, 2024), as well as from game theory formulations via the so-called referential signaling games introduced by Lewis (Lewis, 1969; Huttegger et al., 2010a; Huttegger & Zollman, 2011; Huttegger et al., 2010b; Barrett & Gabriel, 2021; Huttegger et al., 2014).
While, as mentioned above, several works show that IL allows for simplicity- or KC-based regularization (Mordatch & Abbeel, 2018; Li & Bowling, 2019b; Korbak et al., 2020; Ren et al., 2023, 2020; Wärneryd, 1993; Zhou et al., 2022; Li & Bowling, 2019a; Ren et al., 2020), these approaches are often built with language-based, transduction (Raj et al., 2017) or symbolic (Havrylov & Titov, 2017; Mathy & Feldman, 2012; Lazaridou et al., 2018) data in mind, based on their origin in the theory of language optimization (Foerster et al., 2016), attempts for vision have been focussed on notions of compositionality as in points 1 and 2 from our description (Choi et al., 2018; Nassar et al., 2018; Liu et al., 2022a; Devaraj et al., 2020; Latheef et al., 2024; Vani et al., 2021), and have shown limited success on the datasets which we work on (Schott et al., 2022). This latter work is one of the main motivations of this paper.
1.1 Related work and main contributions
As mentioned before, the use of IL for fitting compositional multimodal statistical distributions is underexplored. Related works such as (Chowdhury et al., 2021; Denamganai & Walker, 2020) feature a sender and receiver but do not develop a symbolic bottleneck that implies KC regularization as ours. Further works like (Dessì et al., 2021) focus on other aspects within a somehow related IL framework for image generation but do not tackle the bottleneck structure in depth. As far as we know, the only existing attempt to use emergent language models in image generation tasks is (Latheef et al., 2024), with which our work shares the choice of a comunication bottleneck, used for long-tailed image generation of brain activity patterns. This model, however, does not make use of the IL framework.
Our main contributions are the following:
  • We propose a joint implementation of the Lewis Reconstruction Game and IL framework, which to our best knowledge is the first instance of such framework that takes into account the peculiarities of self-supervised generation tasks.
  • We introduce a novel regularization technique to favor exploration and diversity in encoding, with mathematically justified tunable biases towards compositionality and encoding efficiency.
  • We accompany the above by a rigorous mathematical framework (developed in the appendices), that defines the setting and objectives of compositional generation, and justifies our architecture choices.

### v3

# A Compressive-Expressive Communication Framework for Compositional Representations

@@ Abstract
Compositional generalization—the ability to interpret novel combinations of familiar elements—is a hallmark of human cognition and language. Despite recent advances, deep neural networks still struggle to acquire this property reliably. In this work, we introduce CELEBI (Compressive-Expressive Language Emergence through a discrete Bottleneck and Iterated learning), a novel self-supervised framework for inducing compositionality in learned representations from pre-trained models, through a reconstruction-based communication game between a sender and a receiver. Building on theories of language emergence, we integrate three mechanisms that jointly promote compressibility, expressivity, and efficiency in the emergent language. First, interactive decoding incentivizes intermediate reasoning by requiring the receiver to produce partial reconstructions after each symbol. Second, a reconstruction-based imitation phase, inspired by iterated learning, trains successive generations of agents to imitate reconstructions rather than messages, enforcing a tighter communication bottleneck. Third, pairwise distance maximization regularizes message diversity by encouraging high distances between messages, with formal links to entropy maximization. Our method significantly improves both the efficiency and compositionality of the learned messages on the Shapes3D and MPI3D datasets, surpassing prior discrete communication frameworks in both reconstruction accuracy and topographic similarity. This work provides new theoretical and empirical evidence for the emergence of structured, generalizable communication protocols from simplicity-based inductive biases.
@@ 1 Introduction
In natural languages, compositionality enables humans to communicate an infinite number of ideas using a finite set of elements [9]. This principle enables speakers to flexibly combine known words and structures to convey novel meanings, supporting flexible and generalizable communication across diverse and previously unseen contexts.
Works on the emergence of compositionality in language [26, 27, 29] have argued that two opposing bottlenecks are necessary for the natural selection of compositional languages between generations of speakers. On the one hand, successful communication requires high expressivity in order to usefully describe the world. On the other, models of emergent communication such as the iterated language framework [26] state that natural speakers naturally minimize the complexity of languages through cultural transmission, implying that more simple and compressive languages are more easily passed on to new speakers. These opposite pressures of expressivity and compressibility generate a trade-off, highlighted in several works [27, 16, 2], argued to be optimized by compositional languages. When only one of the two pressures is present, languages tend to degrade: excessive compressibility leads to ambiguous, degenerate languages, while prioritizing expressivity alone results in holistic languages without structure, which tend to be more difficult to teach and learn [41].
Inducing compositional behavior in neural networks remains a major challenge [30, 22, 48, 25]. A growing body of work suggests that a model’s ability to generalize compositionally is highly sensitive to its training conditions [31, 10, 49, 32, 11], including factors such as data distribution, learning objectives, and task design. One promising direction is the study of emergent communication, where discrete languages evolve for coordination between independent neural agents.
In this work, we build on the Lewis reconstruction game framework [47] to study emergent communication in a reconstruction task. A sender encodes an image into a discrete message and a receiver reconstructs the original. Drawing on theories of language evolution [26, 27], we develop a novel framework for inducing compositional communication grounded in simplicity bias.
We introduce CELEBI (Compressive-Expressive Language Emergence through a discrete Bottleneck and Iterated learning). Within this framework, we introduce three mechanisms in the learned communication protocol:
Progressive Decoding, in which the receiver makes reconstruction predictions after each incoming symbol, rather than waiting until the full message is received. This biases the system towards using intermediate reasoning steps, a process shown to support compositional generalization in settings such as chain-of-thought reasoning [56]. While this mechanism does not directly improve reconstruction accuracy, it yields more efficient and structured communication.
Final-state imitation modifies the standard imitation phase in the iterated learning (IL) framework of cultural evolution [27, 28, 29]. Instead of imitating the entire message, the student is trained to reproduce only the final output of the receiver, effectively reducing the information transmitted across generations. This tighter generational bottleneck increases pressure for compressibility in the emergent communication protocol [53], thereby promoting the emergence of compositional structure as a necessary condition for successful transmission. Empirically, this leads to increased compositionality with minimal degradation in reconstruction accuracy.
Pairwise distance maximization as regularization term, i.e. as a further pressure towards increased diversity in the emergent language, that encourages exploration during the imitation phase. This term pushes the student sender to maximize an approximate Hamming distance between messages in a given batch, finding the most "diverse" protocol. We prove that this regularization gives a lower bound for entropy maximization [57], and an upper bound on the contrastive learning loss NT-Xtent [8].